Calculate the freezing and boiling point of a solution containing 5.57 g of sodium chloride (NaCl) and 65.0 g of water. (MM of NaCl = 58.44 g/mol, Kf of water = 1.86˚C/m, Kb of water = 0.52˚C/m)


1
Expert's answer
2021-03-30T07:52:47-0400

Solution:

Calculate solution molality:

Molality of NaCl = Moles of NaCl / kilograms of water

m(NaCl) = (5.57 g NaCl) × (1 mol NaCl / 58.44 g NaCl) × (1 / 0.065 kg water) = 1.466 m NaCl

Molality of NaCl = 1.466 m


When NaCl dissolves, it separates into one Na+ ion and one Cl ion:

NaCl(aq) → Na+(aq) + Cl(aq)

Because it breaks up into two ions, its van 't Hoff factor (i) is 2.


Thus:

BP - boiling point

∆TBP = Kb × m × i

∆TBP = (0.52˚C/m) × (1.466 m) × (2) = 1.52˚C

Water normally boils at 100˚C.

Hence,

BP = 100˚C + 1.52˚C = 101.52˚C

BP = 101.52˚C


FP - freezing point

∆TFP = Kf × m × i

∆TFP = (1.86˚C/m) × (1.466 m) × (2) = 5.45˚C

Water normally freezes at 0˚C.

Hence,

FP = 0˚C - 5.45˚C = -5.45˚C

FP = -5.45˚C


Answer:

The boiling point of a solution is 101.52˚C.

The freezing point of a solution is -5.45˚C.